# Optimal Stopping for Sequential Bayesian Experimental Design

### v1

@@ A.3 Proof of Theorem 3
$=\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}+\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w})\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}.$
Combining the two terms, we have
$\nabla_{w}V^{\pi_{w},\psi_{w}}_{k}(s_{k})$ $=\bm{1}_{s_{k}\notin\mathcal{T}_{k,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$ (43)
$\quad+\bm{1}_{s_{k}\notin\mathcal{T}_{k,w}}\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}.$
@@
$=\bm{1}_{s_{k}\notin\mathcal{T}_{k,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$
$\quad+\bm{1}_{s_{k}\notin\mathcal{T}_{k,w}}\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})$
$\hskip 70.0001pt\cdot\bm{1}_{s_{k+1}\notin\mathcal{T}_{k+1,w}}\nabla_{w}\mu_{k+1,w}(s_{k+1})\nabla_{\xi_{k+1}}Q^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1},\xi_{k+1})\Big|_{\xi_{k+1}=\mu_{k+1,w}(s_{k+1})}\,ds_{k+1}$
$\quad+\bm{1}_{s_{k}\notin\mathcal{T}_{k,w}}\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\bm{1}_{s_{k+1}\notin\mathcal{T}_{k+1,w}}\int_{s_{k+2}}p(s_{k+1}\rightarrow s_{k+2}|\pi_{w},\psi_{w})\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+2}(s_{k+2})\,ds_{k+2}\,ds_{k+1}$
$=\prod_{j=k}^{k}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$
$\quad+\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\prod_{j=k}^{k+1}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k+1,w}(s_{k+1})\nabla_{\xi_{k+1}}Q^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1},\xi_{k+1})\Big|_{\xi_{k+1}=\mu_{k+1,w}(s_{k+1})}\,ds_{k+1}$
$\quad+\int_{s_{k+2}}p(s_{k}\rightarrow s_{k+2}|\pi_{w},\psi_{w})\prod_{j=k}^{k+1}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+2}(s_{k+2})\,ds_{k+2}$
$=\prod_{j=k}^{k}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$
$\quad+\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\prod_{j=k}^{k+1}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k+1,w}(s_{k+1})\nabla_{\xi_{k+1}}Q^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1},\xi_{k+1})\Big|_{\xi_{k+1}=\mu_{k+1,w}(s_{k+1})}\,ds_{k+1}$
$\quad+\int_{s_{k+2}}p(s_{k}\rightarrow s_{k+2}|\pi_{w},\psi_{w})\prod_{j=k}^{k+2}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k+2,w}(s_{k+2})\nabla_{\xi_{k+2}}Q^{\pi_{w},\psi_{w}}_{k+2}(s_{k+2},\xi_{k+2})\Big|_{\xi_{k+2}=\mu_{k+2,w}(s_{k+2})}\,ds_{k+2}$
@@
$\quad+\int_{s_{N}}p(s_{k}\rightarrow s_{N}|\pi_{w},\psi_{w})\prod_{j=k}^{N}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}V^{\pi_{w},\psi_{w}}_{N}(s_{N})\,ds_{N}$
$=\sum_{l=k}^{N-1}\int_{s_{l}}p(s_{k}\rightarrow s_{l}|\pi_{w},\psi_{w})\prod_{j=k}^{l}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\,ds_{l}$
$=\sum_{l=k}^{N-1}\mathbb{E}_{s_{l}|\pi_{w},\psi_{w},s_{k}}\left[\prod_{j=k}^{l}\bm{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\right]\,ds_{l}$ (44)
@@
$=\sum_{l=0}^{N-1}\mathbb{E}_{s_{l}|\pi_{w},\psi_{w},s_{0}}\left[\prod_{j=0}^{l}(1-\bm{1}_{s_{j}\in\mathcal{T}_{j,w}})\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\right]$
$=\sum_{l=0}^{N-1}\mathbb{E}_{s_{l}|\pi_{w},\psi_{w},s_{0}}\left[\bm{1}_{l<\tau_{w}}\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\right].$ (45)
A.4 Implementation details
This section addresses several practical challenges including neural network design for variable-length histories and efficient computation of information-theoretic rewards.
Policy Network Architecture
The policy network $\mu_{w}(k,s_{k})$ takes as input both the current stage index and the experimental history. To handle the temporal nature of sequential experimentation, we design the input representation as follows: For the stage information, we employ one-hot encoding to capture the discrete temporal structure:
$k\qquad\longrightarrow\qquad e_{k+1}=[0,\dots,0,\underbrace{1}_{(k+1)\rm{th}},0,\dots,0]^{T}.$ (46)
for the second component, i.e., the state $s_{k}$ (including both $s_{k,b}$ and $s_{k,p}$), we represent it in a nonparametric manner
which poses a challenge due to its variable length across different stages. We address this by using a fixed-size representation that accommodates the maximum possible history length. For experiments up to stage (N1)(N-1), we allocate a total dimension of $(N-1)(N_{\xi}+N_{y})$ and pad incomplete histories with zeros. The complete input vector for stage $k$ takes the form:
$[\underbrace{e_{k+1}}_{N},\overbrace{\xi_{0}}^{N_{\xi}},\dots,\xi_{k-1},\underbrace{0,\dots,0}_{N_{\xi}(N-1-k)},\overbrace{y_{0}}^{N_{y}},\dots,y_{k-1},\underbrace{0,\dots,0}_{N_{y}(N-1-k)}]^{T}.$ (48)
This representation ensures consistent input dimensionality across all stages while preserving the sequential structure of the experimental process. The total input dimension scales linearly as $N+(N-1)(N_{\xi}+N_{y})$. The output layer is an $N_{\xi}$-dimensional vector representing $\xi_{k}$, and the network architecture can be chosen by the user.
Q-Network Architecture
The Q-network follows a similar input design but includes an additional component for the candidate design $\xi_{k}$, resulting in input dimension $N+(N-1)(N_{\xi}+N_{y})+N_{\xi}$. The network outputs a scalar value representing the expected future return.
Information Gain Computation
The terminal reward calculation requires computing KL divergences between posterior distributions. For problems without analytical posteriors, we approximate these quantities by discretizing the parameter space $\theta$ on a regular grid and estimating posterior densities pointwise.
A.5 Analytical solution to the linear-Gaussian problem
We derive the analytical solution to the linear-Gaussian problem described in section 5.1, under the terminal reward formulation. The conjugate Gaussian structure enables us to obtain closed-form expressions for the optimal stopping and design policies. Due to this conjugacy, both the prior and posterior distributions remain Gaussian throughout the sequential process, with log-density functions given by:
@@ A.5 Analytical solution to the linear-Gaussian problem
We list the optimal utilities for varying number of experiments in Table 1. The optimal stopping stages are highlighted in red.
Table 1: Optimal Utility for varying experimental horizons.
NN 1 2 3 4
$U(\Xi_{N})(c_{k}=0)$ 2.203 2.547 2.749 2.892
$U(\Xi_{N})(c_{k}=-0.5)$ 1.703 1.547 1.249 0.892
$U(\Xi_{N})(c_{k}=-0.25)$ 1.953 2.047 1.999 1.892
A.6 Numerical experiments details
A.6.1 Stopping probability schedule
We use an adaptive sigmoid-like stopping probability schedule, with $p_{\text{stop}}>0.999$ for the final 30 training iterations to guarantee convergence to the true optimal stopping policy. Figure 5 illustrates this schedule over 300 training iterations.
Figure 5: Stopping probability schedule used in the curriculum learning approach over 300 training iterations.
A.6.2 Contaminant source detection problem setup
The contaminant concentration $G$ at spatial location $z=[z_{x},z_{y}]$ and time $t$ follows convection-diffusion PDE:
$\frac{\partial G(z,t;\theta)}{\partial t}$ $=\nabla^{2}G-u(t)\cdot\nabla G+S(z,t;\theta),$
$\qquad\qquad z\in[z_{L},z_{R}]^{2},\quad t>0$ (66)
where $\theta=[\theta_{x},\theta_{y},\theta_{h},\theta_{s}]\in\mathbb{R}^{4}$ parameterizes the Gaussian source function
$S(z,t;\theta)=\frac{\theta_{s}}{2\pi\theta_{h}^{2}}\exp(-\frac{(\theta_{x}-z_{x})^{2}+(\theta_{y}-z_{y})^{2}}{2\theta_{h}^{2}})$ (67)
with location $(\theta_{x},\theta_{y})$, width $\theta_{h}$, and strength $\theta_{s}$. ${u}=[u_{x},u_{y}]\in\mathbb{R}^{2}$ is a time-dependent convection velocity.
A mobile sensor measures noisy concentrations
$y_{k}=G(z=s_{k+1,p},t_{k};\theta)+\epsilon_{k}$ (68)
where $\epsilon_{k}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. The design variable $\xi_{k}$ represents sensor displacement:
$s_{k+1,p}=s_{k,p}+\xi_{k}.$ (69)
We consider $N$ measurement opportunities at times $t_{k}=k\cdot\Delta t$ for $k=0,\ldots,N-1$, with $\Delta t=5.0\times 10^{-4}$. We focus on source localization with $\theta_{h}=0.05$, $\theta_{s}=2$, $u_{x}=u_{y}=10t/0.2$ and prior $\theta_{x},\theta_{y}\sim\mathcal{U}(0,1)$.
Solving the forward model (66) using finite volume is still computationally expensive. We use DNNs to construct surrogate models of $G(z,t_{k};\theta)$ for $k=0,\ldots,N-1$, to accelerate the computation. Each DNN uses a 4-dimensional input layer taking $z$ and $\theta$, 5 hidden layers with 40, 80, 40, 20, and 10 nodes, and a scalar output $G$. A dataset is generated by solving for $G$ on 2000 samples of $\theta$ drawn from its prior distribution. These concentration values are then first restricted to only the domain that is reachable by the vehicle (due to the design constraint), then shuffled across $\theta$ and split $80\%$ for training and $20\%$ for testing.
A.6.3 Additional results
We present distributional analysis and representative episode comparisons to provide deeper insights into the learned policies.
The distributional analysis in Figure 6 for the linear-Gaussian benchmark shows that the trained policy predominantly terminates at stage 3 with designs concentrated near the upper bound ($\xi=3$), consistent with analytical solutions. The stopping stage histogram (left) confirms that the learned policy correctly identifies stage 3 as optimal when experimental costs are absent. The design histogram (right) reveals that optimal designs cluster near the constraint boundary, maximizing the signal-to-noise ratio by selecting the largest feasible design values. This concentration pattern validates that the policy gradient approach successfully learns the theoretically optimal behavior in this benchmark case.
Figure 7 presents representative episodes from both methods in the contaminant source detection problem ($c_{k}=-0.8$). Vanilla policy terminates after two experiments with a total reward of 1.33, while curriculum-trained policy continues for three experiments, achieving a higher total reward of 1.99. The curriculum approach achieves superior expected performance and more stable convergence, which is precisely what the optimization objective seeks to maximize.
(a) Vanilla, $\theta=(0.61,0.73)$, total reward = 1.33 (information gain = 2.93, experimental cost = -1.6)
(b) Curriculum, $\theta=(0.61,0.73)$, total reward = 1.99 (information gain = 4.39, experimental cost = -2.4)
Figure 7: Episode instances for vanilla (upper) and curriculum (lower) policies in the $c_{k}=-0.8$ scenario. Purple star: true source location; red dots: sensor positions; red lines: sensor movements; contours: posterior PDF.
Finally, we explore a more complex cost structure in the contaminant source detection problem with design-dependent experimental costs $c_{k}=-\|\xi_{k}\|^{2}$, where sensor movements to distant locations incur higher penalties. Figure 8 shows training convergence under this quadratic cost structure. Curriculum learning again outperforms the vanilla method, achieving significant higher average rewards and converging to an optimal stopping stage around 3. This scenario demonstrates that the curriculum approach maintains its effectiveness even when cost structures become more complex and spatially dependent, suggesting broader applicability to realistic experimental design problems where costs vary with design choices.
Figure 8: Training convergence of the convection-diffusion problem with design-dependent costs $c_{k}=-\|\xi_{k}\|^{2}$. Left: average reward; right: average stopping stage.

### v2

@@ Appendix A Appendix
$=\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}+\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}.$ (54)
Finally (49) gives us
$\nabla_{w}V^{\pi_{w},\psi_{w}}_{k}(s_{k})$ $=\mathbf{1}_{s_{k}\notin\mathcal{T}_{k,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$ (55)
$\quad+\mathbf{1}_{s_{k}\notin\mathcal{T}_{k,w}}\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}.$
@@
$=\mathbf{1}_{s_{k}\notin\mathcal{T}_{k,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$
$\quad+\mathbf{1}_{s_{k}\notin\mathcal{T}_{k,w}}\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})$
$\qquad\qquad\qquad\cdot\mathbf{1}_{s_{k+1}\notin\mathcal{T}_{k+1,w}}\nabla_{w}\mu_{k+1,w}(s_{k+1})\nabla_{\xi_{k+1}}Q^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1},\xi_{k+1})\Big|_{\xi_{k+1}=\mu_{k+1,w}(s_{k+1})}\,ds_{k+1}$
$\quad+\mathbf{1}_{s_{k}\notin\mathcal{T}_{k,w}}\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\mathbf{1}_{s_{k+1}\notin\mathcal{T}_{k+1,w}}$
$\qquad\qquad\qquad\cdot\int_{s_{k+2}}p(s_{k+1}\rightarrow s_{k+2}|\pi_{w},\psi_{w})\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+2}(s_{k+2})\,ds_{k+2}\,ds_{k+1}$
$=\prod_{j=k}^{k}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$
$\quad+\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\cdot\prod_{j=k}^{k+1}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k+1,w}(s_{k+1})\nabla_{\xi_{k+1}}Q^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1},\xi_{k+1})\Big|_{\xi_{k+1}=\mu_{k+1,w}(s_{k+1})}\,ds_{k+1}$
$\quad+\int_{s_{k+2}}p(s_{k}\rightarrow s_{k+2}|\pi_{w},\psi_{w})\prod_{j=k}^{k+1}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+2}(s_{k+2})\,ds_{k+2}$
$=\prod_{j=k}^{k}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}$
$\quad+\int_{s_{k+1}}p(s_{k}\rightarrow s_{k+1}|\pi_{w},\psi_{w})\cdot\prod_{j=k}^{k+1}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k+1,w}(s_{k+1})\nabla_{\xi_{k+1}}Q^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1},\xi_{k+1})\Big|_{\xi_{k+1}=\mu_{k+1,w}(s_{k+1})}\,ds_{k+1}$
$\quad+\int_{s_{k+2}}p(s_{k}\rightarrow s_{k+2}|\pi_{w},\psi_{w})\cdot\prod_{j=k}^{k+2}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{k+2,w}(s_{k+2})\nabla_{\xi_{k+2}}Q^{\pi_{w},\psi_{w}}_{k+2}(s_{k+2},\xi_{k+2})\Big|_{\xi_{k+2}=\mu_{k+2,w}(s_{k+2})}\,ds_{k+2}$
@@
$\quad+\int_{s_{N}}p(s_{k}\rightarrow s_{N}|\pi_{w},\psi_{w})\prod_{j=k}^{N}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}V^{\pi_{w},\psi_{w}}_{N}(s_{N})\,ds_{N}$
$=\sum_{l=k}^{N-1}\int_{s_{l}}p(s_{k}\rightarrow s_{l}|\pi_{w},\psi_{w})\prod_{j=k}^{l}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\,ds_{l}$
$=\sum_{l=k}^{N-1}\mathbb{E}_{s_{l}|\pi_{w},\psi_{w},s_{k}}\left[\prod_{j=k}^{l}\mathbf{1}_{s_{j}\notin\mathcal{T}_{j,w}}\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\right].$ (56)
@@
$=\sum_{l=0}^{N-1}\mathbb{E}_{s_{l}|\pi_{w},\psi_{w},s_{0}}\left[\prod_{j=0}^{l}(1-\mathbf{1}_{s_{j}\in\mathcal{T}_{j,w}})\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\right]$
$=\sum_{l=0}^{N-1}\mathbb{E}_{s_{l}|\pi_{w},\psi_{w},s_{0}}\left[\mathbf{1}_{l<\tau_{w}}\nabla_{w}\mu_{l,w}(s_{l})\nabla_{\xi_{l}}Q^{\pi_{w},\psi_{w}}_{l}(s_{l},\xi_{l})\Big|_{\xi_{l}=\mu_{l,w}(s_{l})}\right].$ (57)
A.4 Analytical solution to the linear-Gaussian problem
We derive the analytical solution to the linear-Gaussian problem described in Section˜4.1, under the terminal reward formulation. The conjugate Gaussian structure enables us to obtain closed-form expressions for the optimal stopping and design policies. Due to this conjugacy, both the prior and posterior distributions remain Gaussian throughout the sequential process, with log-density functions given by:
@@ A.4 Analytical solution to the linear-Gaussian problem
A.5 Hyperparameters
The hyperparameters used for training are given in Table˜2.
Table 2: Hyperparameter settings for numerical experiments.
Linear Gaussian Nonlinear Source Detection
Policy network architecture $N+(N-1)(N_{\xi}+N_{y})\!\to\!80\!\to\!80\!\to\!N_{\xi}$
Q-network architecture $N+(N-1)(N_{\xi}+N_{y})+N_{\xi}\!\to\!80\!\to\!80\!\to\!1$
Learning rate (policy) $1.5\times 10^{-1}$
Learning rate (critic) $1.0\times 10^{-3}$
Critic batch size $500$
Exploration scale $\sigma_{\text{explore}}$ $1.0$ $0.05$
Exploration scale decay $0.99$
Monte Carlo size $M$ $1000$
A.6 Additional results
While direct baselines for learned stopping in sequential experimental design are scarce in the literature, we construct a principled comparator by formulating stopping as part of the action space within a generic reinforcement learning framework. Specifically, we augment the action with a discrete stop or continue decision modeled by a stochastic policy, coupled with a continuous design policy for experiment selection. Both components are trained under a Soft Actor-Critic (SAC) (Haarnoja et al., 2018) objective, yielding a unified hybrid-control baseline that does not rely on problem-specific stopping heuristics. Figure˜15 presents the comparison on the linear-Gaussian benchmark. Even in this relatively simple setting, the SAC baseline performs substantially worse than our vanilla policy gradient. It fails to identify the optimal stopping stage and consequently achieves a lower total reward.
This behavior reflects a concrete difficulty that arises when applying a standard MDP formulation to stopping problems in sequential BED. The hybrid action space, which combines continuous experimental designs with a discrete stop or continue decision, forces generic policy gradient methods to represent stopping through a stochastic policy, since discrete actions cannot be differentiated directly. This is inconsistent with Theorem˜3.1, which establishes that the optimal stopping policy is deterministic. Our key theoretical contribution is to exploit this structure to bypass the hybrid optimization problem altogether. By showing that the stopping policy is implicitly determined by the design policy, we reduce the joint optimization to a single gradient based update over the continuous design parameters alone, for which efficient pathwise estimators apply. The SAC baseline confirms that this reduction is not merely a modeling convenience. Without exploiting this structure, the mismatch between stochastic stopping and deterministic optimality leads to substantially suboptimal performance even on the simple linear Gaussian benchmark.
Figure 15: Performance comparison between Vanilla PG (our method) and a generic RL baseline (soft actor-critic) on the linear-Gaussian benchmark with zero experimental cost ($c_{k}=0$). Left: average reward; right: average stopping stage.